### v1

@@ Abstract
volume: Volume x, Issue x, 2025
@@ 2 Problem Formulation and Background
The state space ${\mathcal{U}}$ is infinite-dimensional, and the optimization space ${\mathcal{Z}}$ can be finite- or infinite-dimensional. We discretize by letting $\{\eta_{i}\}_{i=1}^{n_{u}}$ be a basis for a finite-dimensional subspace ${\mathcal{U}}_{h}\subset{\mathcal{U}}$ and $\{\varphi_{j}\}_{j=1}^{n_{z}}$ be a basis for a finite-dimensional subspace ${\mathcal{Z}}_{h}\subset{\mathcal{Z}}$ (${\mathcal{Z}}_{h}={\mathcal{Z}}$ for finite-dimensional optimization spaces). The coordinate representations of $u$ and $z$ are denoted as ${\bm{u}}\in{\mathbb{R}}^{n_{u}}$ and ${\bm{z}}\in{\mathbb{R}}^{n_{z}}$. For simplicity, we abuse notation and use $J:{\mathbb{R}}^{n_{u}}\times{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}$, $S:{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}^{n_{u}}$, and $\tilde{S}:{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}^{n_{u}}$ to denote the discretized objective function and solution operators.
@@ 2.1 Post-optimality sensitivity with respect to model discrepancy
Assume that ${\tilde{{\bm{z}}}}$ has been computed, and that $S$ and $\tilde{S}$ have been evaluated $N$ (a small number) times to provide a dataset $\{{\bm{z}}_{\ell},S({\bm{z}}_{\ell})-\tilde{S}({\bm{z}}_{\ell})\}_{\ell=1}^{N}$. Hyper-differential sensitivity analysis with respect to model discrepancy is summarized as the two-step procedure:
@@ 3.2 Prior distribution
$||{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}})||_{{\bm{W}}_{\bm{u}}}^{2}+||\mathbf{L}({\bm{\theta}}){\bm{M}}_{\bm{z}}||_{\mathcal{L}({\bm{M}}_{\bm{z}}{\bm{W}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}},{\bm{W}}_{\bm{u}})}^{2},$ (3.1)
@@
$||{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}})||_{{\bm{W}}_{\bm{u}}}^{2}+||\mathbf{L}({\bm{\theta}}){\bm{M}}_{\bm{z}}||_{\mathcal{L}({\bm{M}}_{\bm{z}}{\bm{W}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}},{\bm{W}}_{\bm{u}})}^{2}={\bm{\theta}}^{T}{\bm{W}}_{\bm{\theta}}{\bm{\theta}},$
@@ 3.3 Trace class covariance operators
For stationary problems, the state space ${\mathcal{U}}$ contains functions $f:\Omega\to{\mathbb{R}}$, where $\Omega\subset{\mathbb{R}}^{s}$, $s=1,2,$ or $3$. In this case, a Laplacian-like operator is given by ${\bm{E}}_{\bm{u}}=\beta_{\bm{u}}{\bm{K}}_{\bm{u}}+{\bm{M}}_{\bm{u}}$, where $\beta_{\bm{u}}>0$ and ${\bm{K}}_{\bm{u}}\in{\mathbb{R}}^{n_{u}\times n_{u}}$ is the stiffness matrix [^1] The $(i,j)$ entry of the stiffness matrix given by the inner product of the gradients of the $i^{th}$ and $j^{th}$ basis functions.. This corresponds to the discretization of the elliptic operator $-\beta_{\bm{u}}\Delta+I$, where $\Delta$ and $I$ are the Laplacian and identity operators, respectively.
@@ 3.4 Computing prior discrepancy samples
where $s_{\bm{z}}=({\bm{z}}-{\tilde{{\bm{z}}}})^{T}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}({\bm{z}}-{\tilde{{\bm{z}}}})$ and $\bm{\omega}_{\bm{u}}^{(1)},\bm{\omega}_{\bm{u}}^{(2)}$ are independent samples from a standard normal distribution on ${\mathbb{R}}^{n_{u}}$. In what follows, we use (3.9) and (3.12) to analyze the prior model discrepancy and guide the specification of its hyper-parameters.
@@ 4.3 State variance hyper-parameter
so that $\mathbb{E}_{\bm{\omega_{\bm{\theta}}}}\left[\|{\delta}({\tilde{{\bm{z}}}},{\bm{T}}_{\bm{\theta}}\bm{\omega}_{\bm{\theta}})\|_{{\bm{M}}_{\bm{u}}}^{2}\right]=\|\bm{d}_{1}\|_{{\bm{M}}_{\bm{u}}}^{2}$.
@@ 4.4 Optimization variable variance hyper-parameter
We leverage evaluations of $\tilde{S}$ to capture characteristic scales that are relevant to provide an estimate of $\gamma^{2}$. For notational simplicity, let $\Delta S(\Delta{\bm{z}})=S({\tilde{{\bm{z}}}}+\Delta{\bm{z}})-S({\tilde{{\bm{z}}}})$ and $\Delta\tilde{S}(\Delta{\bm{z}})=\tilde{S}({\tilde{{\bm{z}}}}+\Delta{\bm{z}})-\tilde{S}({\tilde{{\bm{z}}}})$ and observe that manipulations of (4.3) yields
@@
Figure 2 displays (4.8) as a function of $\beta_{\bm{z}}$ for elliptic operators defined on the unit cube $[0,1]^{s}$ in spatial dimensions $s=1,2,$ and $3$. We observe that (4.8) assumes values within the interval (.$(.2,1)$, that it varies smoothly and monotonically with respect to $\beta_{\bm{z}}$, and that it takes smaller values for higher spatial dimensions. These advantageous properties ensure that we can accurately estimate (4.8) at a modest computational cost.
@@ 4.5 The role of visualization to verify or adjust the hyper-parameters
Recall from (3.12) that ${\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})$ is directly proportional to $\sqrt{\alpha_{\bm{u}}}$ and that its spatial smoothness properties are determined by $\beta_{\bm{u}}$. Inspired by this interpretation of $\alpha_{\bm{u}}$ and $\beta_{\bm{u}}$, we use the maximum absolute value and correlation length of ${\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})$ (which is a function defined on space and/or time) as metrics to project the data $\{{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})\}_{i=1}^{Q}$ onto a 2D scatter plot, which allows for an assessment if the magnitude and smoothness of the prior samples $\{{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})\}_{i=1}^{Q}$ are commensurate to characteristics of $S({\tilde{{\bm{z}}}})-\tilde{S}({\tilde{{\bm{z}}}})$. Visualizing the fields ${\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})$ for the extreme and average cases in the scatter plot reveals if $\alpha_{\bm{u}}$ and $\beta_{\bm{u}}$ have been specified correctly. Generally, $\beta_{\bm{u}}$ should be analyzed first since $\alpha_{\bm{u}}$ depends on it. Furthermore, there are potentially invalid assumptions made in our initialization of $\beta_{\bm{u}}$, whereas the initialization of $\alpha_{\bm{u}}$ did not make assumptions on the data. If the visualizations identify misspecification of the hyper-parameters, the interpretation provided by (3.12) guides how the values should be modified. In particular, if the samples $\{{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})\}_{i=1}^{Q}$ are too smooth (or non-smooth) then $\beta_{\bm{u}}$ should be decreased (or increased) and if their magnitudes are too large (or small) then $\alpha_{\bm{u}}$ should be decreased (or increased).
@@
where $s_{\Delta{\bm{z}}_{k}}=\Delta{\bm{z}}_{k}^{T}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}\Delta{\bm{z}}_{k}=\|{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}\Delta{\bm{z}}\|_{{\bm{M}}_{\bm{z}}}^{2}$. We may interpret $\alpha_{\bm{z}}$ and $\beta_{\bm{z}}$ in (4.10) as controlling the rate of change when ${\delta}$ is perturbed in the direction $\Delta{\bm{z}}_{k}$.
@@ 5.1 Stationary 1D problem
Our first illustration considers robustness of the hyper-parameter initialization algorithms with respect to variability in the data $\{{\bm{z}}_{\ell},\bm{d}_{\ell}\}_{\ell=1}^{N}$. To this end, we vary the number of datapoints $N$, and for each fixed $N$, generate $50$ unique datasets. These $50$ datasets are generated by sampling the source terms $\{{\bm{z}}_{\ell}\}_{\ell=2}^{N}$ from a Gaussian random field (${\bm{z}}_{1}={\tilde{{\bm{z}}}}$ is fixed as the result of solving the low-fidelity optimization problem). The corresponding discrepancy data $\{\bm{d}\}_{\ell=2}^{N}$ is recomputed for each dataset by evaluating the high- and low-fidelity models. Figure 4 displays the resulting hyparameters. Each panel corresponds to a hyper-parameter ($\alpha_{\bm{u}},\beta_{\bm{u}},\alpha_{\bm{z}}$, and $\beta_{\bm{z}}$) and displays the hyper-parameter value as a function of $N$ on the horizontal axis, with variability over the $50$ datasets per $N$ displayed by the range of the vertical lines. For perspective on scales, we average over all datasets and plot horizontal lines to denote ±$\pm 20\%$ variation from the mean hyper-parameter estimate.
@@
Figure 4: The panels from left to right correspond to: $\alpha_{\bm{u}},\beta_{\bm{u}},\alpha_{\bm{z}}$, and $\beta_{\bm{z}}$. The number of datapoints $N$ is varied on the horizontal axis and the range of hyper-parameters generated over all 50 datasets (generated uniquely for each $N$) is depicted by the vertical lines. The horizontal lines denote ±$\pm 20\%$ of the mean hyper-parameter value from all datasets.
@@ 3. $\bullet$
    From fourth row: Decreasing $\beta_{\bm{u}}$ results in discrepancy samples that are less smooth, and have a slightly larger magnitude.
@@ 5. $\bullet$
    From sixth row: Decreasing $\beta_{\bm{z}}$ results in ${\delta}({\bm{z}}^{(2)},{\bm{\theta}}_{i})-{\delta}({\bm{z}}_{1},{\bm{\theta}}_{i})$ having a larger magnitude since its oscillatory characteristic is not damped as significantly.
@@
To demonstrate the importance of the prior in the HDSA-MD framework, we follow the approach presented in [2] to compute the posterior model discrepancy and propagate it through the optimization problem to produce the posterior optimal solution. That is, we compute a probability distribution of the optimal controller, which reflects the uncertainty due to using only two high-fidelity model evaluations. Figure 6 displays the posterior for three sets of prior hyper-parameters. The left panel uses the nominal hyper-parameters corresponding to the second row of Figure 5 and the center and right panels use perturbed hyper-parameters corresponding to the third and sixth rows of Figure 5, respectively. In all three panels, the posterior is compared with the high- and low-fidelity optimal controllers. In practice, the high-fidelity optimal controller cannot be computed, but we include it as a ground truth for comparison here to highlight that an overly restrictive prior (when we divide $\alpha_{\bm{u}}$ by $10$) results in a posterior for which the high-fidelity optimal controller has approximately zero probability.
@@ 5.3 Stationary 2D problem
$-\Delta u+\bm{v}\cdot\nabla u=z\qquad$ on $\text{on }\Omega$
$\nabla u\cdot\bm{n}=0$ on $\text{on }\partial\Omega_{n}$
$u=0$ on $\text{on }\partial\Omega_{d},$
@@
$-\Delta u=z\qquad$ on $\text{on }\Omega$
$\nabla u\cdot\bm{n}=0$ on $\text{on }\partial\Omega_{n}$
$u=0$ on $\text{on }\partial\Omega_{d}.$
@@
It is difficult to interpret the distribution of the samples within each correlation length bin since many samples are projected into the same bin. We augment the scatter plot with a density plot (per bin) to better illustrate the distribution. The correlation length is a measure of $\Delta z_{k}$’s smoothness, so the leftmost samples in the scatter plot correspond to the most non-smooth $\Delta{\bm{z}}_{k}$’s. We select three samples from the leftmost, middle, and rightmost bins. The plots of $\Delta z_{k}$ show increasing smoothness from left to right.
@@
By examining the density centers of each correlation length bin, we observe a relatively large maximum magnitude of ${\delta}(\Delta z_{k},{\bm{\theta}}_{i})$ in the small correlation length bins. This indicates that the discrepancy has a high sensitivity to non-smooth $\Delta z_{k}$’s, which is inconsistent with the physical process of diffusion in the high- and low-fidelity models averaging out such perturbations. The misspecification of $\beta_{\bm{z}}$ was because the data used in the initialization algorithm was a spatially localized source. The empirical correlation length estimation assumes that the data are sampled from a Gaussian random field. However, spatially localized fields are not modeled well via Gaussian random fields, and consequently, the estimate of $\beta_{\bm{z}}$ was too small. The prior is adjusted by increasing $\beta_{\bm{z}}$ and repeating the visualization to confirm that the prior samples have characteristics that are commensurate with what is expected from the physical principles. Additionally, based on knowledge about the high-fidelity model, the samples of $\delta(\Delta z_{k},\theta_{i})$ are about double the magnitude that we expect. We divide $\alpha_{z}$ by $4.0$ to counter this effect.
@@
The second hyper-parameter instance in Figure 8 shows the result after adjusting $\{\alpha_{{\bm{z}}},\beta_{{\bm{z}}}\}$. As in the case before adjustment, we select three samples (from the leftmost, middle, and rightmost correlation length bins) for inspection. The overview plot shows that the overall correlation length is larger as a result of increasing $\beta_{\bm{z}}$, and the magnitude of ${\delta}(\Delta{\bm{z}}_{k},{\bm{\theta}}_{i})$ is smaller as a result of decreasing $\alpha_{\bm{z}}$ (note the range of y-axis is halved from before the adjustment).
@@
The combination of the overview plot and interactive inspection relieved us from the tedious process of trying to assess the quality of hyperparameters by inspecting randomly chosen samples from $\{\Delta{\bm{z}}_{k},{\delta}({\tilde{{\bm{z}}}}+\Delta{\bm{z}}_{k},{\bm{\theta}}_{i})-{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})\}_{i=1,k=1}^{Q,P}$. Specifically, in addition to the quick and intuitive overview, the interactions on the overview plot allow us to quickly inspect outlying samples, such as the leftmost and the rightmost samples, to infer bounds on the samples. If those samples on the boundary of the projected space show unexpected outcomes, we can quickly identify the cause and make necessary adjustments.

### v2

@@ Abstract
volume: Volume x, Issue x, 2025October 8, 2025
@@ 2 Problem Formulation and Background
The state space $\mathcal{U}$} is infinite-dimensional, and the optimization space $\mathcal{Z}$} can be finite- or infinite-dimensional. We discretize by letting $\{\eta_{i}\}_{i=1}^{n_{u}}$ be a basis for a finite-dimensional subspace ${\mathcal{U}}_{h}\subset{\mathcal{U}}$ and $\{\varphi_{j}\}_{j=1}^{n_{z}}$ be a basis for a finite-dimensional subspace ${\mathcal{Z}}_{h}\subset{\mathcal{Z}}$. For finite-dimensional optimization spaces, we have ${\mathcal{Z}}_{h}={\mathcal{Z}}$. Our subsequent analysis focuses on $\mathcal{U}$} and $\mathcal{Z}$} being Sobolev spaces such as $H^{1}$, but we note that $\mathcal{Z}$} can be a finite-dimensional inner product space, such as Euclidean space. Our analysis is valid for such cases, with some simplifications that are highlighted in the article. The coordinate representations of $u$ and $z$ are denoted as ${\bm{u}}\in{\mathbb{R}}^{n_{u}}$ and ${\bm{z}}\in{\mathbb{R}}^{n_{z}}$. For simplicity, we abuse notation and use $J:{\mathbb{R}}^{n_{u}}\times{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}$, $S:{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}^{n_{u}}$, and $\tilde{S}:{\mathbb{R}}^{n_{z}}\to{\mathbb{R}}^{n_{u}}$ to denote the discretized objective function and solution operators.
@@ 2.1 Post-optimality sensitivity with respect to model discrepancy
Assume that $\tilde{{\bm{z}}}$} has been computed, that $S$ and $\tilde{S}$ have been evaluated $N$ (a small number) times to provide a dataset $\{{\bm{z}}_{\ell},S({\bm{z}}_{\ell})-\tilde{S}({\bm{z}}_{\ell})\}_{\ell=1}^{N}$, and that ${\bm{z}}_{1}={\tilde{{\bm{z}}}}$  [^1] The assumption that ${\bm{z}}_{1}={\tilde{{\bm{z}}}}$ may be relaxed. However, it is convenient for subsequent analysis and is typically desirable in practice.. Hyper-differential sensitivity analysis with respect to model discrepancy is summarized as the two-step procedure:
@@ 3.2 Prior distribution
[…]
@@
[…]
@@ 3.3 Trace class covariance operators
For stationary problems, the state space $\mathcal{U}$} contains functions $f:\Omega\to{\mathbb{R}}$, where $\Omega\subset{\mathbb{R}}^{s}$, $s=1,2,$ or $3$. In this case, a Laplacian-like operator is given by ${\bm{E}}_{\bm{u}}=\beta_{\bm{u}}{\bm{K}}_{\bm{u}}+{\bm{M}}_{\bm{u}}$, where $\beta_{\bm{u}}>0$ and ${\bm{K}}_{\bm{u}}\in{\mathbb{R}}^{n_{u}\times n_{u}}$ is the stiffness matrix [^2] The $(i,j)$ entry of the stiffness matrix given by the inner product of the gradients of the $i^{th}$ and $j^{th}$ basis functions.. This corresponds to the discretization of the elliptic operator $\beta_{\bm{u}}\Delta+I$, where $\Delta$ and $I$ are the Laplacian and identity operators, respectively.
@@ 3.4 Computing prior discrepancy samples
[…]
@@ 4.3 State variance hyper-parameter
so that $\mathbb{E}_{\bm{\omega_{\bm{\theta}}}}\left[\|{\delta}({\tilde{{\bm{z}}}},{\bm{T}}_{\bm{\theta}}\bm{\omega}_{\bm{\theta}})\|_{{\bm{M}}_{\bm{u}}}^{2}\right]=\|\bm{d}_{1}\|_{{\bm{M}}_{\bm{u}}}^{2}$. Note that the initialization of $\alpha_{\bm{u}}$ depends only on $\bm{d}_{1}$ and not $\{\bm{d}_{\ell}\}_{\ell=1}^{N}$ since the analysis is based on evaluating the discrepancy at ${\bm{z}}={\tilde{{\bm{z}}}}={\bm{z}}_{1}$.
@@ 4.4 Optimization variable variance hyper-parameter
[…]
@@
Figure 2 displays (4.8) as a function of $\beta_{\bm{z}}$ for elliptic operators defined on the unit cube $0,1]^{s}$ in spatial dimensions $s=1,2,$ and $3$. We observe that (4.8) assumes values within the interval $(.2,1)$, that it varies smoothly and monotonically with respect to $\beta_{\bm{z}}$, and that it takes smaller values for higher spatial dimensions. These advantageous properties ensure that we can accurately estimate (4.8) at a modest computational cost.
@@ 4.5 The role of visualization to verify or adjust the hyper-parameters
[…]
@@
where $s_{\Delta{\bm{z}}_{k}}=\Delta{\bm{z}}_{k}^{T}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}\Delta{\bm{z}}_{k}=\|{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}\Delta{\bm{z}}\|_{{\bm{M}}_{\bm{z}}}^{2}$. We may interpret $\alpha_{\bm{z}}$ and $\beta_{\bm{z}}$ in (4.10) as controlling the rate of change when $\delta$} is perturbed in the direction $\Delta{\bm{z}}_{k}$. Note that $\beta_{\bm{z}}$ is implicit in (4.10) since $\beta_{\bm{z}}$ appears in ${\bm{E}}_{\bm{z}}$, which enters (4.10) via $s_{\Delta{\bm{z}}_{k}}$.
@@ 5.1 Stationary 1D problem
Our first illustration considers robustness of the hyper-parameter initialization algorithms with respect to variability in the data $\{{\bm{z}}_{\ell},\bm{d}_{\ell}\}_{\ell=1}^{N}$. To this end, we vary the number of datapoints $N$, and for each fixed $N$, generate $50$ unique datasets. These $50$ datasets are generated by sampling the source terms $\{{\bm{z}}_{\ell}\}_{\ell=2}^{N}$ from a Gaussian random field (${\bm{z}}_{1}={\tilde{{\bm{z}}}}$ is fixed as the result of solving the low-fidelity optimization problem). The corresponding discrepancy data $\{\bm{d}\}_{\ell=2}^{N}$ is recomputed for each dataset by evaluating the high- and low-fidelity models. Figure 4 displays the resulting hyparameters. Each panel corresponds to a hyper-parameter ($\alpha_{\bm{u}},\beta_{\bm{u}},\alpha_{\bm{z}}$, and $\beta_{\bm{z}}$) and displays the hyper-parameter value as a function of $N$ on the horizontal axis, with variability over the $50$ datasets per $N$ displayed by the range of the vertical lines. For perspective on scales, we average over all datasets and plot horizontal lines to denote $\pm 20\%$ variation from the mean hyper-parameter estimate.
@@
Figure 4: The panels from left to right correspond to: $\alpha_{\bm{u}},\beta_{\bm{u}},\alpha_{\bm{z}}$, and $\beta_{\bm{z}}$. The number of datapoints NN is varied on the horizontal axis and the range of hyper-parameters generated over all 50 datasets (generated uniquely for each NN) is depicted by the vertical lines. The horizontal lines denote $\pm 20\%$ of the mean hyper-parameter value from all datasets.
@@ 3. $\bullet$
    From fourth row: Decreasing $\beta_{\bm{u}}$ results in discrepancy samples that are less smooth, and consequently some samples have extrema with larger magnitudes than in the nominal hyper-parameter case.
@@ 5. $\bullet$
    From sixth row: Decreasing $\beta_{\bm{z}}$ results in ${\delta}({\bm{z}}^{(2)},{\bm{\theta}}_{i})-{\delta}({\bm{z}}_{1},{\bm{\theta}}_{i})$ having a slightly larger magnitude since its oscillatory characteristic is not damped as significantly. This is most easily seen by examining the largest magnitude samples in the leftmost (spatial input =0=0) region of the rightmost panel. For instance, the largest magnitude curve attains a maximum value of approximately $9$, whereas the maximum value is approximately $6$ in the nominal hyper-parameter case.
@@
To demonstrate the importance of the prior in the HDSA-MD framework, we follow the approach presented in [2] to compute the posterior model discrepancy and propagate it through the optimization problem to produce the posterior optimal solution. That is, we compute a probability distribution of the optimal controller, which reflects the uncertainty due to using only two high-fidelity model evaluations. Figure 6 displays the posterior for three sets of prior hyper-parameters. The left panel uses the nominal hyper-parameters corresponding to the second row of Figure 5 and the center and right panels use perturbed hyper-parameters corresponding to the third and sixth rows of Figure 5, respectively. In all three panels, the posterior is compared with the high- and low-fidelity optimal controllers. In practice, the high-fidelity optimal controller cannot be computed, but we include it as a ground truth for comparison here to highlight that an overly restrictive prior (when we divide $\alpha_{\bm{u}}$ by $10$) results in a posterior for which the high-fidelity optimal controller has approximately zero probability. The conclusions drawn from Figure 6 motivate our study of prior model specification. In the subsections that follow, we focus on other facets of the prior model specification with an implicit awareness of the prior’s influence on the posterior optimal solution.
@@ 5.3 Stationary 2D problem
$-\Delta u+\bm{v}\cdot\nabla u=z\qquad$ $\text{on }\Omega$
$\nabla u\cdot\bm{n}=0$ $\text{on }\partial\Omega_{n}$
$u=0$ $\text{on }\partial\Omega_{d},$
@@
$-\Delta u=z\qquad$ $\text{on }\Omega$
$\nabla u\cdot\bm{n}=0$ $\text{on }\partial\Omega_{n}$
$u=0$ $\text{on }\partial\Omega_{d}.$
@@
It is difficult to interpret the distribution of the samples within each correlation length bin since many samples are projected into the same bin. We augment the scatter plot with a density plot (per bin) to better illustrate the distribution. The correlation length is a measure of $\Delta z_{k}$’s smoothness, so the leftmost samples in the scatter plot correspond to the high-frequency (most non-smooth) $\Delta{\bm{z}}_{k}$’s. We select three samples from the leftmost, middle, and rightmost bins. The plots of $\Delta z_{k}$ show decreasing frequency (increasing smoothness) from left to right. Since the model discrepancy prior is smoothing, the high-frequency $\Delta z_{k}$’s lead to lower magnitude changes in the discrepancy $\delta$}. The hyper-parameter $\beta_{\bm{z}}$ controls the relationship between the frequency of $\Delta z_{k}$ and the magnitude of ${\delta}(\Delta z_{k},{\bm{\theta}})$. Taking a large $\beta_{\bm{z}}$ will result in $\delta$} only being sensitive to the lowest frequency $\Delta z_{k}$’s, thus restricting the prior to discrepancies that are nearly constant (as functions of $\bm{z}$}). On the other hand, taking a small $\beta_{\bm{z}}$ will result in $\delta$} exhibiting non-trivial sensitivity to high-frequency $\Delta z_{k}$’s.
By examining the density centers of each correlation length bin, we observe a relatively large maximum magnitude of ${\delta}(\Delta z_{k},{\bm{\theta}}_{i})$ in the small correlation length bins. This observation is based on knowledge of the physical processes in the high- and low-fidelity models. Specifically, the process of diffusion dampens high-frequency sources, and consequently ${\delta}(\Delta z_{k},{\bm{\theta}}_{i})$ should be small for high-frequency $\Delta z_{k}$’s. The magnitude observed in small correlation length bins indicates that $\beta_{\bm{z}}$ is too small.
@@
The misspecification of $\beta_{\bm{z}}$ was because the data used in the initialization algorithm was a spatially localized source. The empirical correlation length estimation assumes that the data is sampled from a Gaussian random field. However, spatially localized fields are not modeled well via Gaussian random fields, and consequently, the estimate of $\beta_{\bm{z}}$ was too small. The prior is adjusted by increasing $\beta_{\bm{z}}$ and repeating the visualization to confirm that the prior samples have characteristics that are commensurate with what is expected from the physical principles.
The samples of $\delta(\Delta z_{k},\theta_{i})$ are about double the magnitude that we expect. Such an assessment is based on physical intuition about what is lacking in the low-fidelity model to prescribe an appropriate discrepancy magnitude. The misspecification of $\alpha_{\bm{z}}$ is due to the many approximations that are used to initialize it. We divide $\alpha_{\bm{z}}$ by $4.0$ to correct the discrepancy magnitude.
@@
The second hyper-parameter instance in Figure 8 shows the result after adjusting $\{\alpha_{{\bm{z}}},\beta_{{\bm{z}}}\}$. As in the case before adjustment, we select three samples (from the leftmost, middle, and rightmost correlation length bins) for inspection. Observe that the magnitude of the lowest-frequency $\Delta z_{k}$ (the rightmost bin) is approximately halved, as a result of changing $\alpha_{\bm{z}}$. Furthermore, the magnitude of the discrepancy decays more rapidly as a function of the correlation length, as a result of changing $\beta_{\bm{z}}$. For instance, around correlation length $0.3$, the median magnitude was previously greater than $15$, but after the update it is less than $5$. Changing $\alpha_{\bm{z}}$ accounts for dividing the magnitude by $2$, while changing $\beta_{\bm{z}}$ accounts for the remaining decrease as high-frequency modes are damped by a larger $\beta_{\bm{z}}$.
@@
The combination of the overview plot and interactive inspection relieved us from the tedious process of trying to assess the quality of hyper-parameters by inspecting randomly chosen samples from $\{\Delta{\bm{z}}_{k},{\delta}({\tilde{{\bm{z}}}}+\Delta{\bm{z}}_{k},{\bm{\theta}}_{i})-{\delta}({\tilde{{\bm{z}}}},{\bm{\theta}}_{i})\}_{i=1,k=1}^{Q,P}$. Specifically, in addition to the quick and intuitive overview, the interactions on the overview plot allow us to quickly inspect outlying samples, such as the leftmost and the rightmost samples, to infer bounds on the samples. If those samples on the boundary of the projected space show unexpected outcomes, we can quickly identify the cause and make necessary adjustments.